# Accurate Small-Signal Modeling of Digitally Controlled Buck Converters with ADC-PWM Synchronization

### v2

@@ I Introduction
To extend the model accuracy near or beyond the switching frequency, the sideband effect must be taken into account to form a multifrequency model[5]. As an extension to SSA, the generalized state-space averaging method (GSSA)[6, 7, 8, 9] introduces Fourier series to account for higher-order harmonics in the time-domain waveforms. Based on the linear time-periodic (LTP) theory[10], the harmonic state-space (HSS)[11, 12, 13] and harmonic transfer function (HTF)[14, 15] methods have been proposed to model frequency coupling effects. HSS employs the harmonic balance principle[16] to represent an LTP system as an infinite-dimensional matrix equation set[5], while HTF captures both direct and cross-frequency interactions. While these multifrequency modeling methods greatly enhance accuracy, they tend to produce high-order models that are computationally intensive. Consequently, computer-based tools are required, and obtaining a simple closed-form solution is often not possible.
@@
Figure 1: PWM modulation schemes with possible synchronized ADC sampling instants (Green) and modulating edges (Red) highlighted. (a) TEM, up-counting sawtooth carrier. (b) LEM, down-counting sawtooth carrier. (c) Symmetrical modulation, triangular carrier.
@@ II Small-Signal Modeling of Digital Buck Converters with Asymmetrical Carriers
To simplify the analysis, the steady-state ADC sampling instant ($kT_{S}$) defines the beginning of each cycle. The block diagram and timing are presented in Fig.2 and Fig.3, where $H_{i}$ is the current sensor gain and $T_{s}$ is the switching period. In practice, $H_{i}$ also includes the ADC gain. The scaled inductor current is denoted as $i_{s}$. The duty-cycle to the inductor current transfer function and the digital PI compensator transfer function[30] are:
@@
Figure 3: Key discrete-domain waveforms of the TEM digital buck converter, the vertical dashed brown line marks the steady-state ADC sampling instant, the vertical dashed green line indicates the ADC sampling instant under perturbation.
The ADC and shadow register are modeled as sample-and-hold blocks. Under the small-signal assumption, the DPWM behaves as an ideal sampler, allowing the duty-cycle perturbation $\hat{d}$ to be treated as an impulse train [29]. When a sampling-hold-sampling sequence operates at the same sampling frequency, it can be equivalently represented by a single sampler followed by a delay equal to the time difference between the two sampling instants (proof provided in Appendix I). By applying this lemma twice, the 3 samplers and 2 ZOH blocks in Fig.2 can be simplified to a single sampler followed by a delay $T_{D}$, without any ZOH. Under the small-signal assumption, $T_{D}$ is defined as the steady-state time between the ADC sampling instant and the moment when the new calculated duty-cycle command takes effect.
Since DPWM-ADC synchronization guarantees that sampling always occurs at the PWM on-interval center, as shown in Fig.4, in each cycle, the ADC sampling instant $t_{smp}$ is related to the on-time perturbation $\hat{t}_{on}$ of the previous cycle:
Figure 4: Sampling instant variation due to DPWM-ADC synchronization. Brown and green vertical dashed lines represent sampling points under steady-state and perturbed, respectively.
@@
(20) agrees with the conclusion of [37] and provides a method to evaluate the analog loop gain from the digital loop gain. Furthermore, (18) resembles the modified Z-transform[39] of $G_{id}H_{i}$ with a sampling delay of $T_{D}$. One can perform partial fraction on $G_{id}H_{i}$ and then utilize the following modified Z-transform identity to evaluate (18):
@@ III Small-Signal Modeling of Digital Buck Converters with Symmetrical Carriers
Figure 6: Key discrete domain waveforms of a digital buck converter with a symmetrical carrier, the horizontal dashed blue line represents the steady-state duty-cycle command, the vertical dashed brown line marks the steady-state ADC sampling instant.
The above analysis for asymmetrical carriers can be extended to the case where the carrier is a symmetric triangular waveform. Fig.6 illustrates the key waveforms of a digital buck converter with a triangular carrier. The ADC samples when $CNTR$ reaches 0, which corresponds to the center of the PWM on-interval. Afterwards, the new duty cycle command is loaded into the shadow register when $CNTR$ reaches $CNTR_{MAX}$. Compared to the asymmetrical case, two important differences arise. First, in the symmetrical case, the ADC sampling instant is always fixed with respect to the carrier and does not vary with any perturbation. Consequently, there will be no $H_{\text{sync}}(z)$. Second, each impulse $\hat{t}_{on}[k]$ now maps to two impulses in $\hat{d}$ that have different delays ($T_{D1}$ and $T_{D2}$), which is similar to the sampled-data COT model in [30].
@@ V Experimental Verification
To further validate the proposed model, a synchronous buck converter prototype was built as shown in Fig. 12. The controller is implemented on a floating-point DSP (TMS320F28379) running at 200 MHz. Both the switching frequency and ADC sampling rate are set to 100 kHz. A high-resolution PWM is employed such that the PWM resolution far exceeds the 12-bit ADC resolution, thus avoiding limit cycles caused by duty-cycle quantization. Because the resolutions of DPWM and ADC are sufficiently high, quantization effects are negligible in the control operation.
@@
To capture parasitics including the ESR and DCR, $G_{id}H_{i}$ is characterized by Bode 100 as:
$G_{id}(s)\times H_{i}=\frac{3091003s+5165486753}{s^{2}+12168.2939s+648181436}\times 0.085$ (27)
@@
The digital converter setup operates under TEM modulation with turn-on centered sampling, with controller gains set as $K_{p}=0.2$ and $K_{i}=0.2$. Fig. 13 shows the measured analog loop gain. The analog loop gain is obtained via Bode 100, using an injection transformer to insert perturbations as a voltage source between the ADC input and the low-side current-sensing resistor. Two input channels of the Bode 100 measure the voltages at both ends of the injection transformer relative to ground.
Fig. 14 presents the comparison of the digital loop gain predicted by the proposed model against measurements obtained via two distinct injection methods. In the first method, akin to that in [36], the DSP samples the perturbation signal from the Bode 100 analyzer, combines it in software with the sampled inductor current, and then outputs both the pre- and post-injection current samples via DACs back to the Bode 100. The second method implements a Software-based Frequency Response Analyzer (SFRA) directly on the DSP: the SFRA injects discrete sinusoidal perturbations in software at selected frequencies and applies a Discrete-Time Fourier Transform (DTFT) to extract the magnitude and phase response at each frequency. This procedure is repeated sequentially over the frequency range to construct the bode plot. Fig. 14 demonstrates that both measurement methods yield equivalent results and closely agree with the model predictions. The discrepancy observed around 4 kHz is primarily attributed to core losses in the inductor during operation, which were not accounted for during the $G_{id}$ characterization.

### v3

@@ I Introduction
To extend the model accuracy near or beyond the switching frequency, the sideband effect must be taken into account to form a multifrequency model[5]. As an extension to SSA, the generalized state-space averaging method (GSSA)[6, 7, 8, 9] introduces Fourier series to account for higher-order harmonics in the time-domain waveforms. Based on the linear time-periodic (LTP) theory[10], the harmonic state-space (HSS)[11, 12] and harmonic transfer function (HTF)[13, 14] methods have been proposed to model frequency coupling effects. HSS employs the harmonic balance principle[15] to represent an LTP system as an infinite-dimensional matrix equation set[5], while HTF captures both direct and cross-frequency interactions. While these multifrequency modeling methods greatly enhance accuracy, they tend to produce high-order models that are computationally intensive. Consequently, computer-based tools are required, and obtaining a simple closed-form solution is often not possible.
@@
Figure 1: PWM modulation schemes with possible synchronized ADC sampling instants (vertical solid, green) and modulating edges (vertical dashed, red) highlighted. (a) TEM, up-counting sawtooth carrier. (b) LEM, down-counting sawtooth carrier. (c) Symmetrical modulation, triangular carrier.
@@ II Small-Signal Modeling of Digital Buck Converters with Asymmetrical Carriers
To simplify the analysis, we define the beginning of each cycle as the steady-state ADC sampling instant ($kT_{S}$). The block diagram and timing are presented in Fig.2 and Fig.3, where $H_{i}$ is the current sensor gain and $T_{s}$ is the switching period. In practice, $H_{i}$ also includes the ADC gain. The scaled inductor current is denoted as $i_{s}$. The duty-cycle to the inductor current transfer function and the digital PI compensator transfer function[29] are:
@@
Figure 3: Key discrete-domain waveforms of a TEM buck converter. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
The ADC and shadow register are modeled as two sets of sample-and-hold blocks. Under the small-signal assumption, the DPWM comparator behaves as another ideal sampler, allowing the duty-cycle perturbation $\hat{d}$ to be treated as an impulse train [28]. When a sampling-hold-sampling sequence operates at the same sampling rate, it can be equivalently represented by a single sampler followed by a delay equal to the time difference between the two sampling instants. By applying this lemma twice, the 3 ideal samplers and 2 ZOH blocks found above can be simplified to a single sampler followed by a delay $T_{D}$, without any ZOH. Under the small-signal assumption, $T_{D}$ is defined as the steady-state time between the ADC sampling instant and the moment when the new calculated duty-cycle command takes effect.
Since DPWM-ADC synchronization guaranties that sampling always occurs at the PWM on-interval center, as shown in Fig.4, in each cycle, the ADC sampling instant $t_{smp}$ is related to the on-time perturbation $\hat{t}_{on}$ of the previous cycle:
Figure 4: Sampling instant variation due to DPWM-ADC synchronization. Dense and sparse vertical dashed lines represent sampling points under steady-state and perturbed, respectively.
@@
(20) agrees with the conclusion of [36] and provides a convenient method to evaluate the analog loop gain from the digital loop gain. Furthermore, (18) resembles the modified Z-transform[38] of $G_{id}H_{i}$ with a sampling delay of $T_{D}$. One can perform partial fraction on $G_{id}H_{i}$ and then utilize the following modified Z-transform identity to evaluate (18):
@@
Figure 6: Key discrete-domain waveforms of a buck converter with a symmetrical carrier. Dense and sparse vertical dashed lines denote the sampling instants at steady state and under perturbation, respectively. The ADC samples at integer multiples of $T_{S}$.
@@ III Small-Signal Modeling of Digital Buck Converters with Symmetrical Carriers
The above analysis for asymmetrical carriers can be extended to the case where the carrier is a symmetric triangular waveform. Fig.6 illustrates the key waveforms of a digital buck converter with a triangular carrier. The ADC samples when $CNTR$ reaches 0, which corresponds to the center of the PWM on-interval. Afterwards, the new duty cycle command is loaded into the shadow register when $CNTR$ reaches $CNTR_{MAX}$. Compared to the asymmetrical case, two important differences arise. First, in the symmetrical case, the ADC sampling instant is always stationary with respect to the carrier and does not vary with any perturbation. Consequently, there will be no $H_{\text{sync}}(z)$. Second, each impulse $\hat{t}_{on}[k]$ now maps to two impulses in $\hat{d}$ that have different delays ($T_{D1}$ and $T_{D2}$), which is similar to the sampled-data COT model in [29].
@@ IV Comparison Between Existing Small-Signal Models and the Proposed Analytical Model
Table IV compares the proposed model with several established modeling approaches, including the sampled-data model in [29], the discrete-time model in [30], and the HTF-based model in [35]. Unlike the existing methods, the proposed model simultaneously incorporates DPWM-ADC synchronization, supports all three DPWM carrier types, and remains computationally inexpensive with low derivation complexity.
@@ V Experimental Verification
To further validate the proposed model, a synchronous buck converter prototype was built as shown in Fig. 12. The controller is implemented on a floating-point DSP (TMS320F28379) running at 200 MHz. Both the switching frequency and ADC sampling rate are set to 100 kHz. A high-resolution PWM is employed such that the PWM resolution far exceeds the 12-bit ADC resolution, thus avoiding limit cycles caused by duty-cycle quantization. Because the resolutions of DPWM and ADC are sufficiently high, quantization effects are negligible in the digital controller operation.
@@
Figure 13: Plant characterization setup. $v_{p}(t)$ is the network analyzer AC output.
To obtain an accurate plant model for validation, the plant transfer function $G_{id}(s)H_{i}$ is experimentally identified using a network analyzer (Bode 100), as illustrated in Fig. 13. The Bode 100 generates an AC perturbation $v_{p}(t)$, which is combined with a DC bias through a resistive network and sampled by the ADC. After offset and scaling in the DSP, this perturbation is superimposed on the steady-state duty-cycle command, producing a perturbed PWM signal $d(t)$. The Bode 100 measures the transfer function from $d(t)$ to $i_{s}(t)$. Ideally, this transfer function is
$G_{id}(s)H_{i}=\frac{i_{s}(s)}{d(s)}=\frac{V_{IN}H_{i}}{V_{D}\times Z_{RLC}(s)},$ (27)
where $V_{D}$ is the amplitude of $d(t)$ and $Z_{RLC}(s)$ is the RLC impedance seen from the switching node $v_{SW}$. Since the open-loop identification is performed on the actual hardware ($V_{IN}=12$, $D=0.256$), all relevant parasitics (including MOSFET on-resistances and current-dependent inductor core losses) are inherently captured. Therefore, the identified transfer function in (28) is used in all subsequent comparisons, so that any discrepancy primarily reflects the accuracy of the proposed small-signal model rather than uncertainties in component values or operating conditions.
$G_{id}(s)H_{i}=12\times\frac{243147.4+429959502}{s^{2}+16837.45s+601451266}\times 0.1$ (28)
In the closed-loop experiments, the converter setup operates under TEM modulation with turn-on centered sampling. PI controller gains are set as $K_{p}=0.2$ and $K_{i}=0.2$. Fig. 14 shows the measured analog loop gain. The analog loop gain is obtained via Bode 100, using an injection transformer to insert perturbations as a voltage source between the ADC input and the low-side current-sensing resistor. Two input channels of the Bode 100 measure the voltages at both ends of the injection transformer relative to the ground.
Fig. 15 presents the comparison of the digital loop gain predicted by the proposed model against measurements obtained via two distinct injection methods. In the first method, akin to that in [35], the DSP samples the perturbation signal from the Bode 100 analyzer, combines it in software with the sampled inductor current, and then outputs both the pre- and post-injection current samples via DACs back to the Bode 100. The second method implements a Software-based Frequency Response Analyzer (SFRA) directly on the DSP: the SFRA injects discrete sinusoidal perturbations in software at selected frequencies and applies a Discrete-Time Fourier Transform (DTFT) to extract the magnitude and phase response at each frequency. This procedure is repeated sequentially over the frequency range to construct the bode plot. Fig. 15 demonstrates that both measurement methods yield equivalent results and closely agree with the model predictions.
@@
VI Conclusions
This paper has presented an exact small-signal model for digitally controlled buck converters operating in FCCM under constant-frequency current-mode control, with explicit consideration of DPWM-ADC synchronization. Based on sampled-data theory, the proposed model rigorously captures all sideband components and provides accurate predictions of both analog and digital loop gains over a wide frequency range, including frequencies beyond the switching and sampling frequencies. This work further shows that, by using the modified $z$-transform, the digital loop gain can be computed in an accurate yet low-complexity closed form, and the analog loop gain can then be directly derived from the digital loop gain, thereby avoiding computationally intensive infinite-series evaluations.
The modeling procedure first considered asymmetrical sawtooth carriers (TEM and LEM) and was then extended to symmetrical triangular carriers. For each case, the sampling-to-PWM actuation delays were summarized in compact expressions that are convenient for control design. For asymmetrical carriers, the proposed analysis, for the first time, identifies and explains an additional feedthrough path introduced by DPWM-ADC synchronization, arising from the fact that the sampling instant moves with duty-cycle perturbations. In contrast, under symmetrical carriers the sampling instant remains fixed with respect to the carrier, no such feedthrough path exists, and the resulting loop dynamics are fundamentally different.
Compared with existing models, the proposed framework, for the first time, systematically treats all practically relevant combinations of carrier types (asymmetrical and symmetrical) and sampling instants in digitally controlled buck converters. In all cases examined, the analytical predictions of loop gains agree well with SIMPLIS simulations in both the analog and digital domains. The experimental results further confirm the accuracy and practical relevance of the proposed model. Moreover, the software-based SFRA yields digital loop-gain measurements that closely match both the proposed model and the network-analyzer-based results, indicating that SFRA can be effectively used to support digital controller design even in the absence of a dedicated network analyzer.
The combination of an analytically tractable model and experimentally validated accuracy makes the proposed approach a useful tool for the design and stability assessment of digitally controlled power converters with synchronized sampling. Future work will focus on extending the methodology to other converter topologies and multi-phase architectures.